OpenStudy (anonymous):

PLEASE HELP I HAVE ONLY TILL 11:59 TO GET THESE QUESTIONS DONE. What is the common difference of a 57-term arithmetic sequence where the first term is -25 and the sum is 17,727? (1 points) Answer 11 11.5 12 12.5 What is the sum of a 14-term arithmetic sequence where the last term is 30 and the common difference is -5? (1 points) Answer 875 812.5 455 422.5

OpenStudy (anonymous):

Then why get behind? Sorry, not trying to be rude, but come on.

OpenStudy (anonymous):

im not behind , im just trying to bring my grade up from a high b to an a.

OpenStudy (anonymous):

i did all the rest of my questions i cant figure out these ones though

OpenStudy (anonymous):

FLVS?

OpenStudy (anonymous):

Are you on FLVS?

OpenStudy (anonymous):

whats that ?

OpenStudy (anonymous):

Lol. Never mind.

OpenStudy (campbell_st):

sum of an arithmetic sequence is \[s _{n} = \frac{n}{2}[2a + (n - 1)d]\] n = 57, a = -25 and s = 17727\[17727 = \frac{57}{2}[2 \times (- 25) + (57 - 1)d]\] then 622 = -50 + 56d solve for d 672 = 56d d = 12

OpenStudy (campbell_st):

question 2 find th 1st term 1st using the formula for a term in an arithmetic series 30 = a + (14 -1) (-5) 30 = a - 65 a = 95 then the sum \[s _{14} = \frac{14}{2}[ 95 + 30]\] the sum is 875